OpenStudy (anonymous):

Tom travels from a to b and at a speed of 90km per hour. It takes him time ‘t’. Alice travels from a to b at a speed of 70km per hour. It takes her time ‘t+ 1 hour How long did it take for Tom to complete his journey?’

OpenStudy (anonymous):

@camerondoherty @Callisto @phi @Abhisar

OpenStudy (anonymous):

im looking

OpenStudy (anonymous):

what are you asking for???the time

OpenStudy (anonymous):

@ryanvarghese12790

OpenStudy (phi):

you start by knowing rate* time = distance next, what do you know about the distance of Tom's trip as compared to Alice's trip ?

OpenStudy (anonymous):

is: \[90t=70(t+1)\] \[20t=70\] Therefore, \[t = \frac{ 70 }{ 20 }\] \[= 3.5 hours\]

OpenStudy (phi):

yes

OpenStudy (anonymous):

cheers for the help ... will give you a medal and fan you....

OpenStudy (phi):

If there is any fun in math, it's figuring out the problem

OpenStudy (anonymous):

ok....lol

OpenStudy (anonymous):

yes its 3.5 hours

OpenStudy (anonymous):

lol...you are bit late... ☺

OpenStudy (anonymous):

oh sry

OpenStudy (anonymous):

no problem ♠

OpenStudy (anonymous):

=]

OpenStudy (anonymous):

♦ ♠ ☺ ↨ ♣ ƒ Ú õ ° ï Ä ı ├Õ

OpenStudy (anonymous):

@camerondoherty how do u get soooo many fans?

OpenStudy (camerondoherty):

I've been on Open Study a while I've made friends, and helped many people. That's how c:

OpenStudy (anonymous):

nice.....